OpenStudy (anonymous):

I just need someone to check this Solve the triangle. A = 52°, b = 10, c = 7

OpenStudy (anonymous):

for a i got 12

OpenStudy (anonymous):

for C i got 43.9

OpenStudy (anonymous):

and for B i got 84.1

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (mathmale):

surely appreciate your having done this much work before asking for help. Congrats. Let's try applying the Law of Cosines to find the length of side a. The Law states:\[a^2=b^2+c^2-2bc*\cos A\]

OpenStudy (mathmale):

Mind calculating ' -2bc*cos A '? Show your work. The reason I ask you to do this is that I get a different result for the length of side a.

OpenStudy (anonymous):

7.9!! thats what i got first for a

OpenStudy (mathmale):

-( b )( c )cos ( A ) = -(10)(7)cos (52 deg) = -70*cos (52 deg) = ??

OpenStudy (mathmale):

Once you have that value, add 10^2 and 7^2 to it. Your result?

OpenStudy (anonymous):

7.9